# A Review of Gravitational Memory and BMS Frame Fixing in Numerical Relativity

### v1

@@ III Literature Review
Gravitational memory effects were first realized in 1974 when Zel’dovich and Polnarev successfully calculated the gravitational radiation that is produced by two objects on flyby, i.e., hyperbolic, trajectories Zel’dovich and Polnarev (1974). [^10] Even earlier, in 1966, Newman and Penrose identified that, near null infinity, the strain of surfaces of constant retarded time will change between spacelike and future timelike infinity Newman and Penrose (1966). By working with Einstein’s equations in linearized GR, they found that, because the stress energy tensor exhibits a net change between early and late times due to the change in the mass distribution of the flyby objects, the strain will also exhibit such a net change. Consequently, because this can also be understood as stemming from a change in the Bondi mass aspect, the effect that they unearthed is what we now call ordinary displacement memory. Later, in 1985, the consequences of their result was elaborated upon by Braginski and Grishchuk who first named this effect the “memory effect” Braginsky and Grishchuk (1985). Then, in 1987, Braginsky and Thorne found a simple equation for the memory for scattering scenarios in terms of the four-momentum of the ingoing and outgoing massive particles Braginsky and Thorne (1987). It says that for a system of $N$ particles, the net change in the gravitational-wave strain between early and late times is
@@
After these early works, it was largely thought that memory effects were understood. This opinion, however, was completely overturned when in 1991 Christodoulou found that gravitational waves themselves will also source a certain type of memory effect, through a subtle, but detectable nonlinear interaction with themselves Christodoulou (1991). [^11] This discovery was also realized by Payne (somewhat indirectly) as well as Blanchet and Damour in Refs. Payne (1983) and Blanchet and Damour (1992). Christodoulou obtained this important result by working with null hypersurface equations and asymptotic limits to obtain an equation that is similar in spirit to Eq. (II.6). In particular, he found that the strain is related to the flux of radiation through each point on the two-sphere. Because of this connection to the energy flux, we now call this effect the null displacement memory. A year later, Thorne identified Christodoulou’s finding as equivalent to that of Ref. Thorne (1992), but with the massive particles being replaced by null gravitons, i.e.,
@@
As for the BMS group, this was realized well before memory effects in 1962 by Bondi, Metzner, van der Burg, and Sachs Bondi et al. (1962); Sachs (1962a). However, the connection between the BMS group and memory was not explicitly stated until Refs. Winicour (2014); Flanagan and Nichols (2015); Strominger and Zhiboedov (2016) in 2014, even though this relationship has been largely understood since, e.g., Refs. Newman and Penrose (1966); Geroch and Winicour (1981); Ashtekar and Streubel (1981); Dray and Streubel (1984); Ludvigsen (1989). What makes this history even more interesting, though, is in Refs. Strominger (2014); Strominger and Zhiboedov (2016) the duality between memory effects and BMS symmetries was extended to be a triangle, since it was found that BMS symmetries and memory effects could also be related to soft theorems Weinberg (1965). This finding created a large stir of interest in theory communities, since soft theorems are inherently useful for studying the $\mathcal{S}$-matrix of a quantum theory, so BMS symmetries are integral to understanding quantum gravity. The “soft limit” means taking a particle’s energy to zero, so soft particles can not be directly measured at colliders. However, memory effects will soon be observed via gravitational waves with current detectors Hübner et al. (2020, 2021); Grant and Nichols (2023), and therefore serve as a natural probe of this exciting realm of physics. This excitement was only further enhanced once more memory effects were unearthed through this connection. In particular, in Refs. Cachazo and Strominger (2014); Pasterski et al. (2016); Nichols (2018), by studying extensions of the BMS group to include “superrotations” de Boer and Solodukhin (2003); Banks (2003); Barnich and Troessaert (2010b, c), i.e., extensions of the Lorentz transformations, two new memory effects were found: the spin memory and the center-of-mass memory. These two effects correspond to the net displacement that two observers with an initial relative velocity will experience due to the passage of transient gravitational radiation. [^12] They can also viewed as the magnetic and electric components of the more general drift memory Grant (2023). Ever since, the field of celestial holography, which aims to firmly establish a holographic dictionary between gravitational scattering in asymptotically flat spacetimes and a certain kind of conformal field theory on the celestial sphere, has reached unprecedented levels of interest Strominger (2017); Pasterski (2019); Raclariu (2021); Pasterski (2021); Pasterski et al. (2021).
@@
One complexity that arose, however, was, with these new NR waveforms that contained memory, it was not exactly obvious how to compare these finite waveforms to post-Newtonian (PN) waveforms that had information about the entire past history of the binary inspiral and thus had a larger prediction for the instantaneous value of the strain Wiseman and Will (1991). This was important because if one wanted to construct a hybridization between these NR waveforms and PN waveforms to build a waveform model from, the hybridization would fail because the two predictions would not agree (see, e.g., Fig. 6). This was resolved in a series a works which established a program now called the BMS frame fixing program Mitman et al. (2021a); Magaña Zertuche et al. (2022); Mitman et al. (2022). Ever since, numerous findings and advancements in gravitational wave physics using these developments have been made Mitman et al. (2023); Yoo et al. (2023); Grant and Mitman (2023); Zhu et al. (2023, 2024).
@@ IV Mathematical Overview
At this point one can then impose Einstein’s equations to obtain evolution equations for the various functions. Specifically, from the $\mathcal{O}(uu,r^{-2})$ and $\mathcal{O}(uA,r^{-2})$ parts [^14] Here the first argument corresponds to the component of the metric tensor that is being examined, while the second argument corresponds to the relevant term in the $1/r$ expansion. of the Einstein tensor one obtains
@@
By contracting these evolution equations with dyads on the two-sphere [^15] See, e.g., Ref. Mitman et al. (2022) for more details. and making use of the Bianchi identities for the Weyl scalars, one can then rewrite Eqs. (17) in terms of spin-weighted functions as
@@
Then, by working with the Bianchi identities, i.e.,
$\dot{\Psi}_{1}$ =ð$=\eth\Psi_{2}+2\sigma\Psi_{3},$ (20a)
$\dot{\Psi}_{2}$ =ð$=\eth\Psi_{3}+\sigma\Psi_{4},$ (20b)
$\mathrm{Im}\left[\Psi_{2}\right]$ $=-\mathrm{Im}\left[\eth^{2}\bar{\sigma}+\sigma\dot{\bar{\sigma}}\right],$ (20c)
$\dot{\Psi}_{3}$ =ð$=-\eth\dot{\bar{\sigma}},$ (20d)
$\dot{\Psi}_{4}$ $=-\ddot{\bar{\sigma}},$ (20e)
reorganizing terms, and integrating with respect to time, one can rewrite Eqs. (19) as
@@
$\mathrm{Im}\left[\eth^{2}\bar{\sigma}\right]$ =ð$=\eth^{-1}\partial_{u}\mathrm{Im}\left[\hat{N}+\mathcal{J}\right],$ (21b)
@@
Eq. (21a) is called the supertranslation conservation law and can be thought of as the conservation law that stems from the supertranslation symmetry of $\mathcal{I}^{+}$; Eq. (21b), however, is not often presented in the literature, in part because it does not have as clear of an interpretation as Eq. (21a) as being the conservation law corresponding to a symmetry contained in the BMS group. Nonetheless, other symmetry groups of future null infinity have also been proposed that contain more symmetries than the BMS group for which Eq. (21b) can be understood as a conservation law de Boer and Solodukhin (2003); Banks (2003); Barnich and Troessaert (2010b, c). Specifically, Eq. (21b) has been referred to as the superrotation conservation law. These other groups are extensions of the BMS group. They are obtained by relaxing the fall off conditions in Eqs. (16a) to be less restrictive, which ends up enabling the existence of other symmetries that extend the usual Lorentz transformations. But, because in this review we restrict our attention to focus on the BMS group, for the remainder of this paper we will not consider these other symmetry groups and instead refer the interested reader to explore these extensions further in Refs. de Boer and Solodukhin (2003); Banks (2003); Barnich and Troessaert (2010b, c).
To provide more motivation as to why, at least to some degree, Eqs. (21) should be viewed as conservation laws, let us first consider what the Bondi mass aspect and the angular momentum aspect represent. In particular, from these aspects one can construct Poincaré charges, i.e., the translation, rotation, and boost charges:
@@
$=\sqrt{\frac{4\pi}{3}}Y_{(1,0)}.$ (24f)
Meanwhile, the quantities $\mathcal{E}$ and $\mathcal{J}$ defined in Eqs. (22) are the usual energy and angular momentum fluxes. So, if one takes Eqs. (21) and instead writes them in terms of spherical harmonics, it can readily be seen that, since ð$\eth^{2}\bar{\sigma}$ has no $\ell=0$ or $\ell=1$ components, these equations are the four and angular momentum conservation laws that correspond to the translation and rotation symmetries. Obtaining such charges can be achieved by following the prescription of Wald and Zoupas Wald and Zoupas (2000) or the derivation of these exact charges by Dray and Streubel Dray and Streubel (1984).
@@
where $p$ and $q$ are real numbers Dray and Streubel (1984). This was pointed out by Dray and Streubel in Ref. Dray and Streubel (1984) and was also later independently realized by Wald and Zoupas in Ref. Wald and Zoupas (2000). From this supermomentum expression, it can be easily shown that if $p+q=1$ then the supermomentum is real and if $p=q$ there is no supermomentum flux in Minkowski space, both of which are nice properties Dray and Streubel (1984). Consequently, the natural choice of supermomentum is the Geroch (G) supermomentum with $p=q=\frac{1}{2}$, i.e.,
@@
At this point, having established the BMS charges, i.e., Eqs. (23) as well as Eq. (25), we can now return to the examination of Eqs. (21) as the BMS conservation laws. In particular, since ð$\eth^{2}\bar{\sigma}$ has no $\ell=0$ or $\ell=1$ components when written in terms of spherical harmonics, we readily find that the $\ell=0$ component of Eq. (21a) is stating energy conservation, while the $\ell=1$ component is stating linear momentum conservation. As for Eq. (21b), the $\ell=0$ component of this equation is trivially zero. But, the $\ell=1$ is stating angular momentum conservation. And, finally, if we now examine the $\ell\geq 2$ components of Eq. (21a), we readily find a statement regarding the conservation of supermomentum. For Eq. (21b), this has sometimes been viewed as a statement regarding the conservation of super angular momentum, but we stress that to do so requires extending the symmetry group of future null infinity to be larger than the BMS group.
@@ V Numerical Code Frameworks
  • extrapolation, which consist of fitting metric data to polynomials in $1/r$ to extract waveform data at future null infinity via Eq. (16); or
@@
The idea of CCE was first theorized in 1996 Bishop et al. (1996, 1997). However, it was not until 2009 when a CCE code was first used on binary black hole merger simulations Reisswig et al. (2010, 2009); Babiuc et al. (2011); Pollney and Reisswig (2011). This version of CCE was run using the finite-difference Pitt Null code Reisswig et al. (2010, 2009); Babiuc et al. (2011); Pollney and Reisswig (2011). In 2014, an improved version of CCE using spectral methods was incorporated into the SpEC code Handmer and Szilagyi (2015); Handmer et al. (2015, 2016). And, finally, in 2020 and 2021 an even more-improved version of CCE that enabled the extraction of the Weyl scalars was developed by Ref. Moxon et al. (2020) and incorporated into the SpECTRE code Deppe et al. (2024) by Ref. Moxon et al. (2023). This version of CCE is the most advanced version and is what will be used throughout this review.
@@ VI Numerical Waveforms
With the pedagogical and mathematical theory behind the BMS group, its corresponding conservation laws, and memory effects presented in Sec. II.6, we now turn to a numerical study of GWs in the context of asymptotics. In particular, to highlight the usefulness of Eqs. (21), we first present Fig. 3 to illustrate how the gravitational wave strain can be easily interpreted in terms of BMS charges and fluxes. In this figure we show four spin weight $-2$ spherical harmonic modes of the gravitational wave strain as well as the contributions to the strain coming from the BMS charges and fluxes in the right-hand sides of Eqs. (21). The system considered is a mass ratio $q=1.22$ binary black hole, whose black hole spins are parallel to the system’s orbital angular momentum and have the dimensionless magnitudes $\chi_{z}^{(1)}=0.33$ and $\chi_{z}^{(2)}=-0.44$. These parameters resemble the most likely parameters of the first gravitational wave detection, GW150914.
@@
The perhaps more interesting panel in Fig. 3, however, is the bottom left panel. In this plot we now show the $(2,0)$ mode of the strain as well as the contributions from both the Bondi mass aspect and the energy flux. As can be seen, instead of the strain being sourced by the Bondi mass aspect, we instead find that there is a large and dominating contribution from the energy flux. Furthermore, there is the unique phenomenon that the strain no longer decays to zero; this is the memory effect! [^18] One also may observe that there is a nonnegligible contribution from the Bondi mass aspect near the peak of the strain. This is because when the black holes are merging, they form an excited remnant black hole that emits gravitational waves in a process called the ringdown. The contribution from the mass aspect is exactly this ringing that occurs as the remnant black hole settles to be in a state of equilibrium, i.e., a Kerr black hole More specifically, this is the term that we associate with the memory. What we mean by this is the following. Formally, gravitational wave memory is a phenomenon that can only be measured between two non-radiative regions of spacetime, e.g., before and after the passage of a burst of gravitational radiation. In practice, however, gravitational wave detectors are not really freely-falling [^19] This is because the test masses in a gravitational wave detector are influenced by actuators that are always working to restore the initial configuration of the detector. and measure the strain in the frequency domain. Thus, to measure memory we need to associate the memory with some nonzero frequency, which is most naturally whatever the frequency of the energy flux is, because it is the source of the memory for these binary systems, as illustrated through the bottom left panel of Fig. 3. So, for the remainder of the paper when we refer to memory what we will really be referring to is the evolution of the memory, as measured through the BMS fluxes.

### v2

@@ III Literature Review
Gravitational memory effects were first realized in 1974 when Zel’dovich and Polnarev successfully calculated the gravitational radiation that is produced by two objects on flyby, i.e., hyperbolic, trajectories Zel’dovich and Polnarev (1974). [^15] Even earlier, in 1966, Newman and Penrose identified that, near null infinity, the strain of surfaces of constant retarded time will change between spacelike and future timelike infinity Newman and Penrose (1966). By working with Einstein’s equations in linearized GR, they found that, because the stress energy tensor exhibits a net change between early and late times due to the change in the mass distribution of the flyby objects, the strain will also exhibit such a net change. Consequently, because this can also be understood as stemming from a change in the Bondi mass aspect, the effect that they unearthed is what we now call ordinary displacement memory. Later, in 1985, the consequences of their result was elaborated upon by Braginski and Grishchuk who first named this effect the “memory effect” Braginsky and Grishchuk (1985). Then, in 1987, Braginsky and Thorne found a simple equation for the memory for scattering scenarios in terms of the four-momentum of the ingoing and outgoing massive particles Braginsky and Thorne (1987). It says that for a system of $N$ particles, the net change in the gravitational wave strain between early and late times is
@@
After these early works, it was largely thought that memory effects were understood. This opinion, however, was completely overturned when in 1991 Christodoulou found that gravitational waves themselves will also source a certain type of memory effect, through a subtle, but detectable nonlinear interaction with themselves Christodoulou (1991). [^16] This discovery was also realized by Payne (somewhat indirectly) as well as Blanchet and Damour in Refs. Payne (1983) and Blanchet and Damour (1992). Christodoulou obtained this important result by working with null hypersurface equations and asymptotic limits to obtain an equation that is similar in spirit to Eq. (II.6). In particular, he found that the strain is related to the flux of radiation through each point on the two-sphere. Because of this connection to the energy flux, we now call this effect the null displacement memory. A year later, in Ref. Thorne (1992) Thorne identified Christodoulou’s finding as equivalent to that of Ref. Braginsky and Thorne (1987), but with the various massive particles being replaced by null radiation, i.e.,
@@
As for the BMS group, this was realized well before memory effects in 1962 by Bondi, Metzner, van der Burg, and Sachs Bondi et al. (1962); Sachs (1962a). However, the connection between the BMS group and memory was not explicitly stated until Refs. Winicour (2014); Flanagan and Nichols (2015); Strominger and Zhiboedov (2016) in 2014, even though this relationship has been largely understood since, e.g., Refs. Newman and Penrose (1966); Geroch and Winicour (1981); Ashtekar and Streubel (1981); Dray and Streubel (1984); Ludvigsen (1989). What makes this history even more interesting, though, is in Refs. Strominger (2014); Strominger and Zhiboedov (2016) the duality between memory effects and BMS symmetries was extended to be a triangle, since it was found that BMS symmetries and memory effects could also be related to soft theorems Weinberg (1965). This finding created a large stir of interest in theory communities, since soft theorems are inherently useful for studying the $\mathcal{S}$-matrix of a quantum theory, so BMS symmetries seem integral to understanding quantum gravity. The “soft limit” means taking a particle’s energy to zero, so soft particles can not be directly measured by particle colliders. However, memory effects will soon be observed through GWs with current detectors Hübner et al. (2020, 2021); Grant and Nichols (2023), and therefore serve as a natural probe of this exciting realm of physics. This excitement was only further enhanced once more memory effects were unearthed through this connection. In particular, in Refs. Cachazo and Strominger (2014); Pasterski et al. (2016); Nichols (2018), by studying extensions of the BMS group to include “superrotations” de Boer and Solodukhin (2003); Banks (2003); Barnich and Troessaert (2010b, c), i.e., extensions of the Lorentz transformations, two new memory effects were found: the spin memory and the center-of-mass memory. These two effects correspond to the net displacement that two observers with an initial relative velocity will experience due to the passage of transient radiation. [^17] They can also be viewed as the magnetic and electric components of the more general drift memory Grant (2023). Ever since, the field of celestial holography, which aims to establish a kind of holographic dictionary between gravitational scattering in asymptotically flat spacetimes and a conformal field theory on the celestial sphere, has reached unprecedented levels of interest Strominger (2017); Pasterski (2019); Raclariu (2021); Pasterski (2021); Pasterski et al. (2021).
@@
One complexity that arose, however, was, with these new NR waveforms that contained memory, it was not exactly obvious how to compare these finite waveforms to post-Newtonian (PN) waveforms that had information about the entire past history of the binary inspiral and thus had a larger prediction for the instantaneous value of the strain Wiseman and Will (1991). This was important because if one wanted to construct a hybridization between these NR waveforms and PN waveforms to build a waveform model from, the hybridization would fail because the two predictions would not agree (see, e.g., Fig. 6). This was resolved in a series of works which established a program now called the BMS frame fixing program Mitman et al. (2021a); Magaña Zertuche et al. (2022); Mitman et al. (2022). Ever since, numerous findings and advancements in gravitational wave physics using these developments have been made Mitman et al. (2023); Yoo et al. (2023); Grant and Mitman (2023); Zhu et al. (2023, 2024).
@@ IV Mathematical Overview
At this point one can then impose Einstein’s equations to obtain evolution equations for the various functions. Specifically, from the $\mathcal{O}(uu,r^{-2})$ and $\mathcal{O}(uA,r^{-2})$ parts [^19] Here the first argument corresponds to the component of the metric tensor that is being examined, while the second argument corresponds to the relevant term in the $1/r$ expansion. of the Einstein tensor one obtains Flanagan and Nichols (2017)
@@
By contracting these evolution equations with dyads on the two-sphere [^20] See, e.g., Ref. Mitman et al. (2020) for more details. and making use of the Bianchi identities for the Weyl scalars, i.e.,
$\dot{\Psi}_{0}$ =ð$=\eth\Psi_{1}+3\sigma\Psi_{2},$ (20a)
$\dot{\Psi}_{1}$ =ð$=\eth\Psi_{2}+2\sigma\Psi_{3},$ (20b)
$\dot{\Psi}_{2}$ =ð$=\eth\Psi_{3}+\sigma\Psi_{4},$ (20c)
$\mathrm{Im}\left[\Psi_{2}\right]$ $=-\mathrm{Im}\left[\eth^{2}\bar{\sigma}+\sigma\dot{\bar{\sigma}}\right],$ (20d)
$\Psi_{3}$ =ð$=-\eth\dot{\bar{\sigma}},$ (20e)
$\Psi_{4}$ $=-\ddot{\bar{\sigma}},$ (20f)
one can then rewrite Eqs. (18) rather simply in terms of spin-weighted functions as
@@
Then, by reorganizing terms (for Eq. (22a)) or following the derivation in Ref. Mitman et al. (2020) (for Eq. (22b)), one can rewrite Eqs. (21) with the shear on the left-hand side as
@@
$\mathrm{Im}\left[\eth^{2}\bar{\sigma}\right]$ $=\partial_{u}\mathrm{Im}\left[\eth^{-1}\left(\hat{N}+\mathcal{J}\right)\right],$ (22b)
@@
In Eqs. (23), $\mathcal{E}$ can be thought of as the total energy flux measured at $\mathcal{I}^{+}$ and $\mathcal{J}$ as the angular momentum flux. Eq. (22a) is called the supertranslation conservation law and can be thought of as the conservation law that stems from the supertranslation symmetry of $\mathcal{I}^{+}$; Eq. (22b), however, is not often presented in the literature, in part because it does not have as clear of an interpretation as Eq. (22a) as being the conservation law corresponding to a symmetry contained in the BMS group. Nonetheless, other symmetry groups of future null infinity have also been proposed that contain more symmetries than the BMS group for which Eq. (22b) can be understood as a conservation law de Boer and Solodukhin (2003); Banks (2003); Barnich and Troessaert (2010b, c). Specifically, Eq. (22b) has been referred to as the superrotation conservation law. These other groups are extensions of the BMS group. They are obtained by relaxing the fall off conditions in Eqs. (17) to be less restrictive, which ends up enabling the existence of other symmetries that extend the usual Lorentz transformations. But, because in this review we restrict our attention to focus on the BMS group, for the remainder of this paper we will not consider these other symmetry groups and instead refer the interested reader to explore these extensions further in Refs. de Boer and Solodukhin (2003); Banks (2003); Barnich and Troessaert (2010b, c).
To provide more motivation as to why, at least to some degree, Eqs. (22) should be viewed as conservation laws, let us first consider what the Bondi mass aspect and the angular momentum aspect represent. [^21] Note that the word aspect here corresponds to the fact that the mass and angular momentum aspects can be thought of as angle-dependent generalizations of mass and angular momentum. In particular, from these aspects one can construct Poincaré charges, i.e., the translation, rotation, and boost charges Gómez López and Quiroga (2017); Dray and Streubel (1984); Dray (1985); Streubel (1978)
@@
$=\sqrt{\frac{4\pi}{3}}Y_{(1,0)},$ (25f)
with each component being a spin-weight 0 function. Meanwhile, the quantities $\mathcal{E}$ and $\mathcal{J}$ defined in Eqs. (23) are the usual energy and angular momentum fluxes. So, if one takes Eqs. (22) and instead writes them in terms of spherical harmonics, it can readily be seen that, since ð$\eth^{2}\bar{\sigma}$ has no $\ell=0$ or $\ell=1$ components, these equations are the four and angular momentum conservation laws that correspond to the translation and rotation symmetries. Obtaining such charges can be achieved by following the prescription of Wald and Zoupas Wald and Zoupas (2000) or the derivation of these exact charges by Dray and Streubel Dray and Streubel (1984).
@@
where $p$ and $q$ are real numbers Dray and Streubel (1984). This was pointed out by Dray and Streubel in Ref. Dray and Streubel (1984) and was also later independently realized by Wald and Zoupas in Ref. Wald and Zoupas (2000). From this supermomentum expression, it can be easily shown that if $p+q=1$ then the supermomentum is real and if $p=q$ there is no supermomentum flux in Minkowski space, both of which are nice properties Dray and Streubel (1984). Consequently, the natural choice of supermomentum is the Geroch (G) supermomentum with $p=q=\frac{1}{2}$ Geroch (1977); Dray and Streubel (1984)
@@
At this point, having established the BMS charges, i.e., Eqs. (24) as well as Eq. (26), we can now return to the examination of Eqs. (22) as the BMS conservation laws. In particular, since ð$\eth^{2}\bar{\sigma}$ has no $\ell=0$ or $\ell=1$ components when written in terms of spherical harmonics, we readily find that the $\ell=0$ component of Eq. (22a) is stating energy conservation, while the $\ell=1$ component is stating linear momentum conservation. As for Eq. (22b), the $\ell=0$ component of this equation is trivially zero. But, the $\ell=1$ is stating angular momentum conservation. And, finally, if we now examine the $\ell\geq 2$ components of Eq. (22a), we readily find a statement regarding the conservation of supermomentum. For Eq. (22b), this has sometimes been viewed as a statement regarding the conservation of super angular momentum, but we stress that to do so requires extending the symmetry group of future null infinity to be larger than the BMS group Compère et al. (2020).
@@ V Numerical Code Frameworks
  • extrapolation, which consists of fitting metric data ($\gamma_{AB}$ in Eq. (17d)) to polynomials in $1/r$ to extract waveform data at future null infinity Boyle et al. (2019); or
@@
The primary output of these extraction methods is the gravitational wave strain $h$, which, with respect to the shear tensor $C_{AB}$ in Eq. (17d), is defined via
$h(u,\theta,\phi)$ $\equiv\bar{q}_{A}\bar{q}_{B}C_{AB}$
$=\sum\limits_{\ell\geq 2,|m|\leq\ell}h_{(\ell,m)}(u)\phantom{}_{-2}Y_{(\ell,m)}(\theta,\phi)$ (29)
and is a spin-weight $-2$ function often decomposed into spin-weight $-2$ spherical harmonics $\phantom{}{}_{-2}Y_{(\ell,m)}$ with the complex coefficients $h_{(\ell,m)}(u)$. In Eq. (V), $q_{A}$ is the dyad $q_{A}\equiv-\left(1,i\sin(\theta)\right)/\sqrt{2}$ (see, e.g., Ref. Mitman et al. (2020)).
@@
The idea of CCE was first theorized in 1996 Bishop et al. (1996, 1997). However, it was not until 2009 when a CCE code was first used on a binary black hole simulation Reisswig et al. (2010, 2009); Babiuc et al. (2011); Pollney and Reisswig (2011). This version of CCE was run using the finite-difference Pitt Null code Reisswig et al. (2010, 2009); Babiuc et al. (2011); Pollney and Reisswig (2011). Later, in 2014, an improved version of CCE using spectral methods was incorporated into the SpEC code Handmer and Szilagyi (2015); Handmer et al. (2015, 2016). And, finally, throughout 2020 and 2021 an even more-improved version of CCE that could extract the Weyl scalars was developed by Ref. Moxon et al. (2020) and incorporated into the SpECTRE code Deppe et al. (2024) by Ref. Moxon et al. (2023). This version of CCE is the most advanced version and is what will be used throughout this review.
@@ VI Numerical Waveforms
With the pedagogical and mathematical theory behind the BMS group, its corresponding conservation laws, and memory effects presented in earlier sections, we now turn to a numerical study of GWs in the context of asymptotics. In particular, to highlight the usefulness of Eqs. (22), we first present Fig. 3 to illustrate how the gravitational wave strain can be easily interpreted in terms of BMS charges and fluxes. In this figure we show four spin weight $-2$ spherical harmonic modes of the gravitational wave strain as well as the contributions to the strain coming from the BMS charges and fluxes in the right-hand sides of Eqs. (22). The system considered is a mass ratio $q=1.22$ binary black hole, whose black hole spins are parallel to the system’s orbital angular momentum and have the dimensionless magnitudes $\chi_{z}^{(1)}=0.33$ and $\chi_{z}^{(2)}=-0.44$. These parameters resemble the most likely parameters of the first gravitational wave detection, GW150914 Abbott et al. (2016).
@@
The perhaps more interesting panel in Fig. 3, however, is the bottom left panel. In this plot we now show the $(2,0)$ mode of the strain as well as the contributions from both the Bondi mass aspect and the energy flux. As can be seen, instead of the strain being sourced by the Bondi mass aspect, we instead find that there is a large and dominating contribution from the energy flux. Furthermore, there is the unique phenomenon that the strain no longer decays to zero; this is the memory effect! [^24] One also may observe that there is a nonnegligible contribution from the Bondi mass aspect near the peak of the strain. This is because when the black holes are merging, they form an excited remnant black hole that emits gravitational waves in a process called the ringdown. The contribution from the mass aspect is exactly this ringing that occurs as the remnant black hole settles to be in a state of equilibrium, i.e., a Kerr black hole More specifically, this is the term that we associate with the memory. What we mean by this is the following. Formally, gravitational wave memory is a phenomenon that can only be measured between two non-radiative regions of spacetime, e.g., before and after the passage of a burst of gravitational radiation. In practice, however, gravitational wave detectors are not really freely-falling [^25] This is because the test masses in a gravitational wave detector are influenced by actuators that are always working to restore the initial configuration of the detector. and analyze the strain in the frequency domain. Thus, to measure memory we need to associate the memory with some nonzero frequency, which is most naturally whatever the frequency of the energy flux is, because it is the source of the memory for these binary systems, as illustrated through the bottom left panel of Fig. 3. So, for the remainder of the paper when we refer to memory what we will really be referring to is the evolution of the memory, as measured through the BMS fluxes.